Shift Summary for eve shift, 01/20/2000


Subject: Shift Summary for eve shift, 01/20/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:34:32 CST


              Shift Summary for eve shift, 01/20/2000

On Shift: Hogan, J. Shields

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of KAMI data (special runs not included): 6.7
Hours of special runs (laser scans, muon runs...): 1.5

Accounting of lost time:

two accesses & a tape change (concurrent w. the access)
========================================================================
Runs ended this shift:

Run 15643 - BEAM_MUON
Run 15644 - BEAM_MUON
Run 15645 _ Pedestal
Run 15646 - KAMI_Ke3

Tapes brought to FCC this shift:
  (including tapes for online split)

QKE201-218, QKB169-170

========================================================================
Accesses during shift: 2

one to check AGEL alignment, & one to reset mainframe 10 and etot

========================================================================
Overview of shift:

80 physicists fail to find h-dibaryon

AGEL muon runs, relatively smooth KAMI_Ke3 datataking, a couple of
brief accesses, PLUS much much discussion about what to do with the
next 3 shifts before beam goes away.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15643 15:30 as new shift arrives, the villagers gather in
                        the control room to discuss which special runs
                        should be done on friday, since KAMI is ahead
                        of schedule

              15:45 realizing that the hope of resolving the issue in
                      "10 or 15 minutes" is futile, the villagers
                      gather up their torches and pitchforks and
                      regroup at New Muon.

              15:50 ADC profiles show that there is a slight bias,
                      indicating a AGEL misalignment. An access is
                      requested and run is stopped.

              16:17 Access is made to check AGEL alignment

15644 1 16:22 begin new muon run

              16:30 ADC profiles now look nicer.

              16:36 end muon run

15645 16:45 begin pedestal run, request 8e12 SEM (which will
                      continue until midnight or when the AGEL study
                      is done). NM2SEED1 from beam to reduce dose
                      rates near target area.

15646 16:55 begin KAMI_Ke3 run

              17:30 Water chiller for FASTBUS tripped. Etot temp
                      high. Duty mechanic paged.

              17:45 Etot trips, but does not affect this run. Duty
                      mechanic successfully resets chiller. End of
                      crisis. JH will reset Etot at the next
                      convienient access.

              20:45 Arun shows up with new Di-baryon plots, which
                      turn out to be disappointingly inconclusive.
                      ( See KAMI logbook for plots).

              21:28 mainframe 10 dies. Shift crew decides to stop
                      run, change tapes, and let JH make his access
                      to reset etot.

              22:00 JH returns from access and resets mf10.

15647 22:10 begin new KAMI_Ke3 run

  

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:34:34 CST